OpenStudy (anonymous):

Maria's penny bank is 1/5 full. After she adds 360 pennies, it is 7/10 full. How many pennies can Maria's bank hold?

OpenStudy (anonymous):

write the equation in variables using given information.

OpenStudy (anonymous):

what is \[\frac{7}{10}-\frac{1}{5}\]?

OpenStudy (anonymous):

@satellite73 1/2

OpenStudy (anonymous):

yes. 1/2

OpenStudy (anonymous):

half of 360 is 180

OpenStudy (anonymous):

360 pennies is 1/2 of bank. So 360*2=720 pennies can be hold in bank